OpenStudy (anonymous):

find the absolute maxima, minima of sin(2x)-2cos(x) on {-pi/2 , pi/2}

OpenStudy (anonymous):

\[\sin(2x)-2\cos(x)\]

OpenStudy (anonymous):

i took the derivative

OpenStudy (anonymous):

\[\cos(2x)+\sin(x)=0 is what i get\]

OpenStudy (anonymous):

someone in my class solved it by doing this: cos(2x)=sin(x) ...

OpenStudy (anonymous):

why and how?

OpenStudy (ankit042):

You can do with derivative expand cos2x = cos^2 -sin^2 +sinx 1-2sin^ 2+sin = 0 (sinx-1)(2sinx +1) = 0 sinx = 1/2 or sinx =1

OpenStudy (ankit042):

sorry sinx = -1/2 and sinx =1

OpenStudy (anonymous):

so you used a trig identity?

OpenStudy (ankit042):

yes cos2x = cosx^2 -sinx^2 and sin^ +cos^ =1

OpenStudy (anonymous):

ok, i see what you did. you found 2 critical points right?

OpenStudy (ankit042):

Yes one for max value and other for min

OpenStudy (anonymous):

alright, and are there other ways to solve this. like using: cos(2x)=sin(x)

OpenStudy (ankit042):

umm I am not sure from where you got cos(2x)=sin(x)

OpenStudy (anonymous):

that's where i'm confused, this guy solved it somehow like this

OpenStudy (anonymous):

and i can barely read his writing

OpenStudy (ankit042):

from our equation we will get cos(2x) = -sin(x)

OpenStudy (anonymous):

exactly..

OpenStudy (ankit042):

are you sure his method gave him right solution.

OpenStudy (anonymous):

\[Let f(x)=\sin 2x-2\cos x\] \[f \prime \left( x \right)=2\cos 2x+2\sin x \] f'(x)=0 gives \[2\cos 2x+2\sin x=0, or \cos 2x+\sin x=0\] \[1-2\sin ^{2}x+\sin x=0\] \[2\sin ^{2}x-\sin x-1=0\] \[2\sin ^{2}x-2\sin x+\sin x-1=0\] \[2 \sin x\left( \sin x-1 \right)+1\left( \sin x+1 \right)=0\] \[\left( \sin x+1 \right)\left( 2\sin x+1 \right)=0\] \[\because given interval is \left( \frac{ -\pi }{ 2 },\frac{ \pi }{ 2 } \right)\] \[either sinx=-1=-\sin \frac{ \pi }{ 2 }=\sin \left( \frac{ -\pi }{ 2 } \right),x=\frac{ -\pi }{ 2 }\] \[\sin x=\frac{ -1 }{ 2 }=-\sin \frac{ \pi }{ 6 }=\sin \frac{ -\pi }{ 6 },x=\frac{ -\pi }{6 }\] \[substitute x=\frac{ -\pi }{6 },\frac{ -\pi }{2 },\frac{ \pi }{ 2 }\] in f(x) and find the absolute max. & min. value

OpenStudy (anonymous):

hm, alright, thanks. i'll just do it this way

OpenStudy (anonymous):

this is the correct way.

OpenStudy (anonymous):

@surjithayer (sinx+1)(2sinx+1)=0 this part of your solution is wrong

OpenStudy (anonymous):

there should be a - 1 instead of +1

OpenStudy (anonymous):

ok so, how do i find the absolute maxima and minima i plugged in -pi/2 and pi/2 in my original function and get 0...what do i do

OpenStudy (anonymous):

i'm guessing the minima is -pi/2 and maxima is pi/2 in the original function

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

but i was told to find the absolute extrema, not local

OpenStudy (anonymous):

i'm so confused, maybe i should take break

OpenStudy (ankit042):

@rowa tell me what's the problem?

OpenStudy (anonymous):

i'm trying to graph it, i don't know how to know if it's going up or down. and if i have a domain of the derived function equal to 0 that is the same domain as the endpoint of my closed interval, can i use it..blablabla

OpenStudy (anonymous):